OpenStudy (sf_r):

Medals** What is the exact area of a circle with a radius of 31 cm? *The answers may come in handy, I've been tryna solve this, but I haven't got the exact answer. A. 15.5 PI cm2 B. 31 PI cm2 C. 961 PI cm2 D. 240.25 PI cm2

OpenStudy (okkaaayyyy):

it will be b! the area is radius multiplied by pi.

OpenStudy (mathstudent55):

Area of circle = square of radius times pi \(\Large A_{circle} = \pi r^2\) \(\Large A_{circle} = \pi (31~cm)^2\) What is \(31^2\) ?